# On Generalized Characters Whose Values on Nonidentity Elements Are Sums of at most Two Roots of Unity

### v1

@@ Abstract.
A character of a finite group having degree $n$ takes values which may be expressed as sums of $n$ or fewer roots of unity. In this note, we prove a result which describes the irreducible constituents of generalized characters on abelian groups whose values on nonidentity elements are expressible as sums of two or fewer roots of unity. In §4, we apply our main result to obtain a result about the connected components of a certain subgraph of the prime graphs for groups admitting such generalized characters.
@@ 1. Introduction
Permutation groups acting with relatively large minimal degree have received extensive study for their applications to Riemann surfaces, and much is known about the structure of such groups (cf. [3, Theorem 1.1]). In particular, the Zassenhaus groups provide us with many irreducible examples of the characters considered in case (iii) of Theorem 2, since in this case, the group acts doubly transitively, and so, the constituent $\chi$ in the permutation character $1+\chi$ is irreducible. Further, examples of that described in case (ii) in the above corollary can occur and with nontrivial center. For instance, consider the transitive action of $\operatorname{SL}(2,3)$ on the cosets of one of its Sylow 3-subgroups. One verifies that under this action, the constituent $\chi$ of the permutation character $1+\chi$ takes values $\pm 1$ off the identity, and so, no element fixes exactly one point.
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3})$.
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}-\lambda_{3})$.
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})$.
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3}-\lambda_{4})$.
@@
Given any generalized character whose values on nonidentity elements are sums of few roots of unity, we would intuitively expect that the generalized character would not differ too much from the regular character, and this result shows this to usually be the case under these circumstances, provided that the prime divisors of the group order are sufficiently large. We also mention that we have examples of each of the four outlier cases, showing that we do indeed need to make exceptions for these cases. For instance, in the case of (iv), take $G$ to be the cyclic group of order 12 and let the $\lambda_{i}$ be the four linear characters (in any order) whose kernels contain the subgroup of order 3. The cyclic group of order 15 has a counterexample of the type in case (i) and of the type in case (ii) as well. An example of type (iii) is also found in the cyclic group of order 9. Further, the requirement that $|G|\leq 21$ cannot be sharpened. By employing GAP, we found that one can choose five distinct linear characters of the cyclic group of order 21 whose values on the nonidentity elements are sums of two roots of unity. Similarly, there exists a sum of seven distinct irreducible characters of the cyclic group of order 15 satisfying this property.
@@
If $G$ is a group and $\Gamma(G)$ has $k$ connected components, there exists $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ whose set of values on $G^{\#}$ is precisely $\{a_{i}\}_{i=0}^{k-1}$ where the $a_{i}$ are distinct.
@@
Assume that $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ and that the values of $\chi$ on $G^{\#}$ each lie in the set $\{-2,-1,0,1,2\}$. Further assume that $\chi$ takes at least 3 of these five values. Is the prime graph of $G$ guaranteed to be disconnected?
Keeping the above notation, if the set of values of $\chi$ on $G^{\#}$ is precisely the set $\{-1,0,1\}$, then Theorem 2 implies the answer is yes.
@@ 2. Preliminaries
We call a vanishing sum of roots of unity minimal if the terms of the sum cannot be partitioned into proper subsums which are themselves vanishing. The above lemma essentially says that the only minimal vanishing sums consisting of $p$-power roots of unity are the “rotations” of $\sum_{i=0}^{p-1}\zeta_{p}^{i}$ by some root of unity $\epsilon$. A classical result of de Bruijn in [1] provides us with an analogous result for vanishing sums when two primes are involved. Alternatively, see [5, Theorem 3.3] for a modern ring-theoretic proof.
@@ 3. Proofs of Main Result
(1) $\chi=\sum_{\lambda\in\operatorname{Irr}(G)}a_{\lambda}\lambda$
@@
(5) $a_{0}x^{2}+a_{1}(x+1)^{2}+a_{2}(x+2)^{2}\leq\dfrac{4(a_{0}+a_{1}+a_{2}-1)}{a_{0}+a_{1}+a_{3}}<4$
or
(6) $(a_{0}+a_{1}+a_{2})x^{2}+2(a_{1}+2a_{2})x+\bigg(a_{1}+4a_{2}-4\bigg)<0.$
The inequality
(7) $a_{0}x^{2}+a_{1}(x+1)^{2}+a_{2}(x+2)^{2}\leq\dfrac{4(a_{0}+a_{1}+a_{2}-1)}{a_{0}+a_{1}+a_{2}},$
when viewed as a quadratic inequality in $x$, has discriminant
(8) $4(a_{1}+2a_{2})^{2}-4(a_{0}+a_{1}+a_{2})\bigg(a_{1}+4a_{2}-\dfrac{4(a_{0}+a_{1}+a_{2}-1)}{a_{0}+a_{1}+a_{2}}\bigg)$
@@
(9) $4<\dfrac{a_{0}a_{1}+a_{1}a_{2}+4a_{0}a_{2}}{a_{0}+a_{1}+a_{2}-1}.$
If either $a_{0}$ or $a_{2}$ equals one, then one verifies that this inequality forces the other to be at most 3. Notice that if $a_{0}=1$ and $a_{2}=3$ (or vice versa), equality actually holds. Likewise, if $a_{0}$ and $a_{2}$ are both at least 2, then the above inequality holds, so this case cannot hold either.
On the other hand, if $a_{2}=0$, then the above inequality becomes
(10) $4(a_{0}+a_{1}-1)-a_{0}a_{1}<0$
@@
First, assume that both $\delta_{i}$’s are both zero, so that $\lambda_{1}(g)+\lambda_{2}(g)+\lambda_{3}(g)-\lambda_{4}(g)=0$ is a minimal vanishing sum of weight 4. If $g$ is chosen to have odd order, then the first three terms have odd order, an impossibility for a vanishing sum of this weight by Lemma 10. Thus, $2\in\pi(G)$.
Now, assume $\delta_{1}=\pm 1$ and $\delta_{2}=0$ (or vice versa). This gives us again a vanishing sum of weight five, and it follows as in Case 3 that either $3\in\pi(G)$ or $5\in\pi(G)$.
Now, assume that the $\delta_{i}$’s are nonzero, so that we have a vanishing sum of weight 6. Notice that in this case, we have equality in (4), and it follows by the triangle inequality that $\delta_{1}\epsilon_{1}=\delta_{2}\epsilon_{2}$. In particular, this vanishing sum cannot be minimal since the terms of a minimal vanishing sum of this weight are each distinct by Lemma 11. If the sum can be partitioned into two minimal vanishing sums of weight 3, then it follows as in Case 3 that at least one of the $\lambda_{i}(g)$’s has order divisible by $3$, and so, $3\in\pi(G)$. On the other hand, we may have that the sum can be partitioned into three minimal vanishing sums of weight 2. If $G$ has odd order, then we can choose $g\in G$ such that the first three $\lambda_{i}$ take mutually distinct values on $g$ by Lemma 12. If $\lambda_{4}(g)=\lambda_{i}(g)$ for some $i\neq 4$ (say $i=3$ for instance), then this implies $\lambda_{1}(g)+\lambda_{2}(g)-\pm 2\epsilon_{1}=0$, and thus, $\lambda_{1}(g)=\lambda_{2}(g)$, a contradiction. On the other hand, if $\lambda_{4}(g)\neq\lambda_{i}(g)$ for each $i$, then $\lambda_{4}(g)$ cancels with one of the $\delta_{i}\epsilon_{i}$ terms. This gives rise to a vanishing sum of weight 4 having three terms of odd order, which is again impossible.
@@
We have seen that $a_{0}\geq 5$ implies $a_{2}=0$ and vice versa. Thus, we may assume without loss that $a_{2}=0$ and $a_{1}\geq 5$. We have seen that this implies $|G|\leq 21$. The only groups of order at most 21 and indivisible by either 2, 3, or 5 are the cyclic groups of order 7, 11, 13, 17, and 19. Further analysis of the inequality in (10) shows that $a_{1}=6$ implies $a_{0}\leq 10$, and $a_{1}=7$ implies $a_{0}\leq 8$. Cases with higher values of $a_{1}$ are subsumed by the cases considered here by symmetry. In particular, it follows that in this case, $\chi=a\rho_{G}\pm\sum_{i=1}^{k}\lambda_{i}$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$ and $5\leq k\leq 7$.
@@
We wonder more generally about generalized characters of abelian groups whose values on nonidentity elements are sums of at most $k$ roots of unity. However, the case analysis required here indicates that it may be difficult to obtain a meaningful result this general.
@@ 4. Prime Graphs
We now turn to proving a result on prime graphs as an application of our main theorem. For a group $G$, $\Gamma(G)$ is taken to be the prime graph of $G$.
@@
Let $G$ be a group, and assume $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ takes values on $G^{\#}$ which are sums of at most two roots of unity. Also, let $\Gamma^{\prime\prime}(G)$ denote the induced subgraph of $\Gamma(G)$ on the set $\pi(G)-\{2,3\}$, and let $\Gamma_{i}^{\prime\prime}(G)$ denote the induced subgraph of $\Gamma(G)$ on the vertex set $\pi_{i}$. Then, if $\pi_{i}$ is nonempty, $\Gamma_{i}(G)$ is a union of connected components of $\Gamma^{\prime\prime}(G)$.
@@
The nonexistence of examples in (ii) was verified using GAP. ∎
We at last have our final result on prime graphs which says that 2 is a “cut vertex” of the prime graph under our assumptions, that is, the prime graph becomes disconnected when discarding the vertex 2.
@@
However, if $\chi_{H}$ has the latter two types, this implies $|H|\leq 16$ and so $p=5$, but this is impossible by the above lemma. If $\chi_{H}$ has either of the first two types, this implies that 3 is linked to an element of $\pi_{0}$. If $G$ has another subgroup $K$ of order $3q$ for some prime $q$ such that $\chi_{K}$ has type $(3,0)$ or $(0,3)$, then $p=5$ by our previous lemma. However, these two types cannot occur within $G$ simultaneously since this would imply that $\chi(1)\equiv-2\ (\mathrm{mod}\ 5)$ and $\chi(1)\equiv 2\ (\mathrm{mod}\ 5)$. In any case, 3 can at most link $\pi_{0}$ to either $\pi_{-2}$ or $\pi_{2}$, but not both simultaneously. More importantly, there is no path in $\Gamma^{\prime}(G)$ linking an element of $\pi_{1}\cup\pi_{-1}$ to an element of $\pi_{0}\cup\pi_{2}\cup\pi_{-2}$.

### v2

@@ Abstract.
A character of a finite group having degree $n$ takes values which may be expressed as sums of $n$ or fewer roots of unity. In this note, we prove a result which describes the irreducible constituents of generalized characters on abelian groups whose values on nonidentity elements are expressible as sums of two or fewer roots of unity. In §4, we apply our main result to obtain information about the connectivity of prime graphs for groups admitting such characters.
Key words and phrases:
generalized characters, abelian groups, finite groups, prime graphs
2020 Mathematics Subject Classification:
20C15, 20K01 (Primary) 20B05, 05C25 (Secondary)
@@ 1. Introduction
Permutation groups acting with relatively large minimal degree have received extensive study for their applications to Riemann surfaces, and much is known about the structure of such groups (cf. [4, Theorem 1.1]). In particular, the Zassenhaus groups provide us with many irreducible examples of the characters considered in case (iii) of Theorem 2, since in this case, the group acts doubly transitively, and so, the constituent $\chi$ in the permutation character $1+\chi$ is irreducible. Further, examples of that described in case (ii) in the above can occur and with nontrivial center. For instance, consider the transitive action of $\operatorname{SL}(2,3)$ on the cosets of one of its Sylow 3-subgroups. One verifies that under this action, the constituent $\chi$ of the permutation character $1+\chi$ takes values $\pm 1$ off the identity, and so, no element fixes exactly one point.
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3})$,
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}-\lambda_{3})$,
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3}+\lambda_{4})$, or
@@
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3}-\lambda_{4})$,
@@
Given any generalized character whose values on nonidentity elements are sums of few roots of unity, we would intuitively expect that the generalized character would not differ too much from the regular character, and this result shows this to usually be the case under these circumstances, provided that the prime divisors of the group order are sufficiently large. We also mention that we have examples of each of the four outlier cases, showing that we do indeed need to make exceptions for these cases. For instance, in the case of (iv), take $G$ to be the cyclic group of order 12 and let the $\lambda_{i}$ be the four linear characters (in any order) whose kernels contain the subgroup of order 3. The cyclic group of order 15 has a counterexample of the type in case (i) and of the type in case (ii) as well. An example of type (iii) is also found in the cyclic group of order 9. Further, the requirement that $|G|\leq 21$ cannot be sharpened. By employing GAP [2], we found that one can choose five distinct linear characters of the cyclic group of order 21 whose values on the nonidentity elements are sums of two roots of unity. Similarly, there exists a sum of seven distinct irreducible characters of the cyclic group of order 15 satisfying this property.
@@
If $G$ is a group and $\Gamma(G)$ has $k$ connected components, there exists $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ whose set of values on $G^{\#}$ is precisely $\{a_{i}\}_{i=0}^{k-1}$ where the $a_{i}$ are distinct integers.
@@
Assume that $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ and that the values of $\chi$ on $G^{\#}$ are sums of at most two roots of unity. Further assume that there exists $x,y,z\in G^{\#}$ such that $\chi(x)=0$, $\chi(y)$ is a root of unity, and $\chi(z)$ is a sum of two roots of unity. Is the prime graph of $G$ guaranteed to be disconnected?
Although we are unable to answer this question at the moment, we are however able to show in Theorem 17 that under these circumstances, the graph becomes disconnected when the vertex corresponding to the prime 2 is thrown away.
@@ 2. Preliminaries
We call a vanishing sum of roots of unity minimal if the terms of the sum cannot be partitioned into proper subsums which are themselves vanishing. The above lemma essentially says that the only minimal vanishing sums consisting of $p$-power roots of unity are the “rotations” of $\sum_{i=0}^{p-1}\zeta_{p}^{i}$ by some root of unity $\epsilon$. A classical result of de Bruijn in [1] provides us with an analogous result for vanishing sums when two primes are involved. Alternatively, see [6, Theorem 3.3] for a modern ring-theoretic proof.
@@ 3. Proofs of Main Result
$\chi=\sum_{\lambda\in\operatorname{Irr}(G)}a_{\lambda}\lambda$
@@
(2) $a_{0}x^{2}+a_{1}(x+1)^{2}+a_{2}(x+2)^{2}\leq\dfrac{4(a_{0}+a_{1}+a_{2}-1)}{a_{0}+a_{1}+a_{3}}.$
The above inequality will be fundamental. When viewing this as a quadratic inequality in $x$, the discriminant is given by
$4(a_{1}+2a_{2})^{2}-4(a_{0}+a_{1}+a_{2})\bigg(a_{1}+4a_{2}-\dfrac{4(a_{0}+a_{1}+a_{2}-1)}{a_{0}+a_{1}+a_{2}}\bigg).$
@@
$4<\dfrac{a_{0}a_{1}+a_{1}a_{2}+4a_{0}a_{2}}{a_{0}+a_{1}+a_{2}-1}.$
If either $a_{0}$ or $a_{2}$ equals one, then one verifies that this inequality forces the other to be at most 3. Notice that if $a_{0}=1$ and $a_{2}=3$ (or vice versa), equality actually holds. Likewise, if $a_{0}$ and $a_{2}$ are both at least 2, then the above inequality holds, so this case cannot hold either. On the other hand, if $a_{2}=0$, then the above inequality becomes
(3) $4(a_{0}+a_{1}-1)-a_{0}a_{1}<0.$
@@
Notice that in this case, we have equality in (2), and it follows by the triangle inequality that $\delta_{1}\epsilon_{1}=\delta_{2}\epsilon_{2}$. In particular, this vanishing sum cannot be minimal since the terms of a minimal vanishing sum of this weight are each distinct by Lemma 11. If the sum can be partitioned into two minimal vanishing sums of weight 3, then it follows as in Case 3 that at least one of the $\lambda_{i}(g)$’s has order divisible by $3$, and so, $3\in\pi(G)$. On the other hand, we may have that the sum can be partitioned into three minimal vanishing sums of weight 2. If $G$ has odd order, then we can choose $g\in G$ such that the first three $\lambda_{i}$ take mutually distinct values on $g$ by Lemma 12. If $\lambda_{4}(g)=\lambda_{i}(g)$ for some $i\neq 4$ (say $i=3$ for instance), then this implies $\lambda_{1}(g)+\lambda_{2}(g)-\pm 2\epsilon_{1}=0$, and thus, $\lambda_{1}(g)=\lambda_{2}(g)$, a contradiction. On the other hand, if $\lambda_{4}(g)\neq\lambda_{i}(g)$ for each $i$, then $\lambda_{4}(g)$ cancels with one of the $\delta_{i}\epsilon_{i}$ terms. This gives rise to a vanishing sum of weight 4 having three terms of odd order, which is impossible.
@@
We have seen that $a_{0}\geq 5$ implies $a_{2}=0$ and vice versa. Thus, we may assume without loss that $a_{2}=0$ and $a_{1}\geq 5$. We have seen that this implies $|G|\leq 21$. The only groups of order at most 21 and indivisible by either 2, 3, or 5 are the cyclic groups of order 7, 11, 13, 17, and 19. Further analysis of the inequality in (3) shows that $a_{1}=6$ implies $a_{0}\leq 10$ and that $a_{1}=7$ implies $a_{0}\leq 8$. Cases with higher values of $a_{1}$ are subsumed by the cases considered here by symmetry. In particular, it follows that in this case, $\chi=a\rho_{G}\pm\sum_{i=1}^{k}\lambda_{i}$ for distinct $\lambda_{i}\in\operatorname{Irr}(G)$ and $5\leq k\leq 7$.
@@
We have actually proven a few other facts within this proof, which we state now for future reference.
Corollary 13.
Assume that $G$ is an abelian group having a generalized character $\chi$ whose values on $G^{\#}$ are sums of at most two roots of unity, and assume that $\chi=a\rho_{G}\pm\sum_{i=1}^{k}\lambda_{i}$ for $a\in\mathbb{Z}$ and distinct $\lambda_{i}\in\operatorname{Irr}(G)$. If $k=6$, then $|G|\leq 16$. If $k=7$, then $|G|\leq 15$.
Also, assume that either of the following holds:
  • (i)
    $\chi=a\rho_{G}\pm(\lambda_{1}+\lambda_{2}+\lambda_{3}-\lambda_{4})$ for $a\in\mathbb{Z}$ and distinct $\lambda_{i}\in\operatorname{Irr}(G)$,
  • (ii)
    $|G|=21$ and $\chi=a\rho_{G}\pm\sum_{i=1}^{5}\lambda_{i}$ for $a\in\mathbb{Z}$ and distinct $\lambda_{i}\in\operatorname{Irr}(G)$,
  • (iii)
    $|G|=16$ and $\chi=a\rho_{G}\pm\sum_{i=1}^{6}\lambda_{i}$ for $a\in\mathbb{Z}$ and distinct $\lambda_{i}\in\operatorname{Irr}(G)$, or
  • (ii)
    $|G|=15$ and $\chi=a\rho_{G}\pm\sum_{i=1}^{7}\lambda_{i}$ for $a\in\mathbb{Z}$ and distinct $\lambda_{i}\in\operatorname{Irr}(G)$.
Then for each $x\in G^{\#}$, we have that $\chi(x)=2\epsilon$ where $\epsilon$ is a root of unity.
We wonder more generally about generalized characters of abelian groups whose values on nonidentity elements are sums of at most $k$ roots of unity. However, the case analysis required here indicates that it may be difficult to obtain a meaningful result in this level of generality.
@@ 4. Prime Graphs
We now turn to proving a result on prime graphs as an application of our main theorem. For a (not necessarily abelian) group $G$, $\Gamma(G)$ is taken to be the prime graph of $G$.
@@
Let $G$ be a group, and assume $\chi\in\mathbb{Z}[\operatorname{Irr}(G)]$ takes values on $G^{\#}$ which are sums of at most two roots of unity. Also, let $\Gamma^{\prime\prime}(G)$ denote the induced subgraph of $\Gamma(G)$ on the set $\pi(G)-\{2,3\}$, and let $\Gamma_{i}(G)$ denote the induced subgraph of $\Gamma(G)$ on the vertex set $\pi_{i}$. Then, if $\pi_{i}$ is nonempty, $\Gamma_{i}(G)$ is a union of connected components of $\Gamma^{\prime\prime}(G)$.
@@
The nonexistence of examples in (ii) was verified using GAP [2]. ∎
We at last have our final result on prime graphs which says that 2 is often a “cut vertex” of the prime graph under our assumptions, that is, the prime graph becomes disconnected when discarding the vertex 2.
@@
However, if $\chi_{H}$ has the latter two types, this implies $|H|\leq 16$ by Corollary 13, and so $p=5$, but this is impossible by the above lemma. If $\chi_{H}$ has either of the first two types, this implies that 3 is linked to an element of $\pi_{0}$. If $G$ has another subgroup $K$ of order $3q$ for some prime $q$ such that $\chi_{K}$ has type $(3,0)$ or $(0,3)$, then $p=5$ by our previous lemma. However, these two types cannot occur within $G$ simultaneously since this would imply that $\chi(1)\equiv-2\ (\mathrm{mod}\ 5)$ and $\chi(1)\equiv 2\ (\mathrm{mod}\ 5)$. In any case, 3 can at most link $\pi_{0}$ to either $\pi_{-2}$ or $\pi_{2}$, but not both simultaneously. More importantly, there is no path in $\Gamma^{\prime}(G)$ linking an element of $\pi_{1}\cup\pi_{-1}$ to an element of $\pi_{0}\cup\pi_{2}\cup\pi_{-2}$.